

********************************************************************
UARS WEEKLY STATUS REPORT - 16 JAN 2004
********************************************************************
***********
SOLSTICE
***********
During this week SOLSTICE continued to operate in good health gathering
solar data on all available orbital tracking periods. On average, we
received 57% of each day's data amounting to about 539 minutes of solar
observing time per day.
***********
SUSIM
***********
SUSIM successfully gathered all of its daily mid (1.1 nm)
resolution, Mg II, and high resolution monitoring scans
during this past week for periods of UARS real-time telemetry
or onboard tape recording.
***********
PEM
***********
Level 2 PEM HEPSA data files are being reprocessed for submission to the
DAAC. This data has been designated as version 2; however, there is no format
difference from the version 1 data format. Version 2 HEPSA data is being
generated because it was discovered that there exists processing errors within
some of the HEPSA version 1 data files. Thus, if you are using any HEPSA
version 1 data files from the DAAC, please replace them with the version 2
data files. Currently, HEPSA data between 1991274 and 2003149 has been
reprocessed and archived at the GSFC DAAC.
Level 2 PEM MEPS data file documentation is complete and submitted to
the DAAC. MEPS data files are version 1. Processed MEPS data files at the
DAAC archive are between 1991274 and 1992399. MEPS data files between
1992340 and 1992349 are waiting for DAAC collection.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered. NOAA has changed the binary structure of the NMC data and
PEM is now receiving two data streams, one in the original VAX format and the
other from their new binary format.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 27 December 2003 (Orbit 67255, GMT Day 361) through
02 January 2004 (Orbit 67359, GMT Day 002).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
NONE THIS REPORT PERIOD
The beta angle ranged between the angles of 65.1 degrees to 79.0
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 0 degrees on 20 January 2004 (DOY 020).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.31 to 0.63 degrees C. Additional
battery performance data for each battery on 26 December and 02
January 2004 are provided in Appendix D.
The clock error ranged from 17.5 to -09.4 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
PEM = 01
FOT = 07
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 04
Four (4) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
No operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
Another DSN test with the spacecraft was performed on 30 December
2003. During this test, data was received at the POCC from the
spacecraft via JPL. The UARS TAC showed lock on the data, but
unfortunately no data was decommed on the Eclipse system. This means
that either the problem lies in the TAC desktop configuration or in
the Eclipse. Analysis will be done next week to determine where the
problem lies.
The FOT continued to investigate the large attitude errors seen on the
simulator during each run of the day-in-the-life (DITL). The power
issues that caused the attitude to diverge were corrected but
subsequent runs still showed the spacecraft (on the simulator) lose
control. More analysis is needed to determine the reason behind the
large errors. This problem needs to be resolved so that the DITL
testing can continue.
MEETINGS
The FOT met with the UARS NOPE and test team at JPL on 29 December
2003 to discuss the subsequent day's test.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 20 January 2004 (DOY 020).
This maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 02 January 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
67265 361/1737 03-337 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
67265 361/1737 03-338 PEM ZEPS/NEPS TO DAY-ONLY
67267 361/2003 03-339 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
67268 361/2142 03-340 OBC/CLK CLK ADJ FR -20.6 TO 41.2 MSEC/DAY
67311 364/1814 03-341 CADH CONFIG TRANSMITTER-2 FOR DSN OMNI
67311 364/1911 03-342 CADH CONFIG TRANSMITTER-2 FOR TDRS OMNI
67311 364/1911 03-343 OBC/CLK CLK ADJ FR 41.2 TO 20.6 MSEC/DAY
67338 001/1254 04-001 OBC/CLK CLK ADJ FR 20.6 TO 41.2 MSEC/DAY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 02 January 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 02 January 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
67265 361/1741 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
67267 361/2010 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
67268 361/2207 CLK ADJ FR -20.6 TO 41.2 MSEC/DAY
67311 364/1911 CLK ADJ FR 41.2 TO 20.6 MSEC/DAY
67338 001/1256 CLK ADJ FR 20.6 TO 41.2 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 02 January 2004
26 December 2003 - GMT Day 360 Beta = 63.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.67/-4.62 +00.0/+00.0 07.2 0.00 0.0
2 +11.2/-22.4 +3.87/+2.91 +20.4/-15.2 26.4 1.02 17.5
3 +11.2/-33.6 +2.91/+1.97 +20.0/-14.8 26.4 1.01 17.0
02 January 2004 - GMT Day 002 Beta = 79.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -1.80/-3.05 +00.0/+00.0 07.2 0.00 0.0
2 +22.4/-22.4 +2.28/+1.65 +12.8/-14.0 27.7 1.07 11.5
3 +33.6/-78.4 +1.65/+1.34 +12.8/-13.6 27.5 1.07 11.0
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period
APPENDIX E
UARS ANOMALY REPORTS
Week ending 02 January 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
03-060 FOT 67255 361/0101 N/A CONFIGS LOCKED UP
03-055 FOT 67272 362/0433 30832 Generic Late Acquisition
03-055 FOT 67278 362/1412 30832 Generic Late Acquisition
04-001 FOT 67343 001/2126 30832 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
03-056 FOT 66992 343/1259 34488 Data Drops
03-057 ATC 67044 347/0100 N/A S/W Stored Cmd Error
03-058 HRDI 67059 348/0101 N/A Not accepting loads
03-059 FOT 67098 350/1515 34541 TDRSS East MA Problem
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
03-055 FOT See List Dec 03 30832 Generic Late Acquisition
03-060 FOT 67255 361/0101 N/A CONFIGS LOCKED UP